How much energy is required to heat 36.0 g H2O from a liquid at 65°C to a gas at 115°C? The following physical data may be useful.

ΔHvap = 40.7 kJ/mol
Cliq = 4.18 J/g°/sup>C
Cgas = 2.01 J/g°/sup>C
Csol = 2.09 J/g°/sup>C
Tmelting = 0°/sup>C
Tboiling = 100°/sup>C
Answer #1

Q1 = 36.0 x 3.18 ( 100 - 65)=5267 J
moles water = 36.0 / 18 g/mol = 2
Q2 = 2 mol x 40.7 kJ=81.4 kJ = 81400 J
Q3 = 36.0 x 2.01 ( 115 - 100)= 1085 J)

Q1 + Q2 + Q3 = 87752 J